. Find the value of Z if the normal curve area between 0 and Z (positive) is 0.4332.
1
Expert's answer
2020-05-20T17:53:47-0400

Solve: Normal curve area between 0 and Z is equal to "\int_0^Z" "\frac{1}{\sqrt{2\pi}}e^\frac{-x^2}{2}dx=\Phi(Z)-\Phi(0)=\Phi(Z)," where "\Phi" is Laplas function.

We have: "\int_0^Z \frac{1}{\sqrt{2\pi}}e^\frac{-x^2}{2}dx" "= 0.4332=\Phi(Z)". Using the table of values of Laplas function we get: Z = 1.5


Answer: Z=1.5